Use the information below for the following 3 questions A recent survey of 80 IU students and 60 Ball State students found that the percent of IU students who are part of Greek life is 18%, while the percent Ball State students who are part of Greek life is 15%. Find the 95% confidence interval for the difference between the proportion of IU students who are part of Greek life compared to Ball State students, i.e. LaTeX: D=p_{IU}-p_{Ball\:State} D = p I U − p B a l l S t a t e . [.0222 ; .0378] [-.0740 ; .1340] [-.1340 ; .0740] [-.0935 ; .1535] Flag this Question Question 13 1 pts Suppose you want to test for whether Greek participation is greater at IU than at Ball State. Which of the following would be the hypothesis test? LaTeX: \left(D=p_{IU}-p_{Ball\:State}\right) ( D = p I U − p B a l l S t a t e ) LaTeX: H_0 \colon D\ge0 ; H_A \colon D<0 H 0 : D ≥ 0 ; H A : D < 0 H 0 : D ≥ 0 ; H A : D < 0 LaTeX: H_0 \colon D=0 ; H_A \colon D\ne0 H 0 : D = 0 ; H A : D ≠ 0 H 0 : D = 0 ; H A : D ≠ 0 LaTeX: H_0 \colon D\ne0 ; H_A \colon D=0 H 0 : D ≠ 0 ; H A : D = 0 H 0 : D ≠ 0 ; H A : D = 0 LaTeX: H_0 \colon D\le0 ; H_A \colon D>0 H 0 : D ≤ 0 ; H A : D > 0 H 0 : D ≤ 0 ; H A : D > 0 Flag this Question Question 14 1 pts What is the test statistic for the difference between IU and Ball State Greek life participation, using the sample proportions found above? Recall:LaTeX: D=p_{IU}-p_{Ball\:State} D = p I U − p B a l l S t a t e (Round intermediate calculations to the third decimal) 0.471 -7.394 1.348 2.209
Answer #1

Solution

Let p1 and p2 represent the population proportion of IU students who are part of Greek life and Ball State students respectively.

Q1 : Confidence Interval

100(1 - α) % Confidence Interval for (p1 - p2) is:

[(p1hat – p2hat) ± MoE, …………………………………………………….. (1)

where

MoE = (Zα/2)√[phat(1 – phat){(1/n1) + (1/n2)}] ……………………………….. (2)

with

p1hat = (x/n1),

p2hat = (y/n2),

phat = (n1 p1hat + n2 p2hat)/( n1 + n2),

Zα/2 is the upper (α /2)% point of N(0, 1),

n1 and n2 being the two sample sizes.

So, 95% confidence interval for the difference between the proportion of IU students who are part of Greek life compared to Ball State students is: [- 0.0949, 0.1549) Fourth Option Answer 1

Details of calculations

n1

80

n2

60

x

14.4

y

9

p1hat

0.18

p2hat

0.15

phat

0.16714286

α

0.05

Zα/2

1.95996398

MoE

0.12488798

LB

-0.094888

UB

0.154888

Q2 : Test Hypotheses

Hypotheses:

Null H0 : p1 = p2   Vs HA : p1 > p2Answer 2

Q3 Test Statistic

Z = (p1hat – p2hat)/√[phat(1 - phat){(1/n1) + (1/n2)} = 0.471 First Option Answer 3

where p1hat and p2hat are sample proportions,

n1, n2 are sample sizes and

phat = {(n1 x p1hat) + (n2 x p2hat)}/(n1 + n2).

Details of calculations

n1

80

n2

60

x

14.4

y

9

p1hat

0.18

p2hat

0.15

phat

0.167142857

Zcal =

0.470813262

DONE